OpenStudy (anonymous):

A line passes through the point (-3, -3) and has a slope of 1/2. What is the equation of the line?

OpenStudy (anonymous):

y-(-3)=1/2(x-(-3)) y+3=1/2 x +3/2 y=(1/2)x - 3/2